OpenStudy (anonymous):

how long to the nearest year will it take an investment in Germany to double its value if the interest is compounded every six months

OpenStudy (tkhunny):

There is not enough information to answer this question as presented. An interest rate is required.

OpenStudy (kropot72):

\[A=P(1+r)^{t}\] When interest is compounded annually the above equation applies, where A is the amount after t years, P is the principal, r is the annual interest rate expressed as a decimal and t is the time in years. When the interest is compounded every six months the equation becomes: \[A=P(1+\frac{r}{2})^{t}\] where t is the number of six monthly intervals. For the principal to double we get: \[\frac{A}{P}=2=(1+\frac{r}{2})^{t}\] Taking logs of both sides: \[t \times \ln (1+\frac{r}{2})=\ln 2\] \[t=\frac{\ln (1+\frac{r}{2})}{\ln 2}\] Whenthe value of t is found it needs to be haved to find the number of years.